# Test-negative designs with various reasons for testing: statistical bias and solution

### v2

@@ Introduction
$VE(\boldsymbol{x})=1-\frac{\Pr(I=1|V=1,\boldsymbol{X}=\boldsymbol{x},R=\textrm{Disease-unrelated reasons},T=1)}{\Pr(I=1|V=0,\boldsymbol{X}=\boldsymbol{x},R=\textrm{Disease-unrelated reasons},T=1)}$
As a result, the relative risk obtained from tested individuals with disease-unrelated reasons targets $VE(\boldsymbol{x})$, covers both healthcare-seekers and non-seekers, and directly assesses vaccine effects on infections without any assumption but effects on symptoms.
Case contact tracing is a common practice of monitoring the spread of diseases in an outbreak setting by testing all contacts of known confirmed cases. Because of the mandatory nature of such contact tracing, testing is unrelated to healthcare-seeking behavior or vaccination status, but it can be related to covariates (e.g., occupation), and contacts of known cases may be more likely to suffer infection and symptoms ($R\rightarrow I$ and $R\rightarrow S$ direct effects in Figure 2(b)). Given these assumptions, we define
$VE_{cct}(\boldsymbol{x})$ $=1-\frac{\Pr(I=1|V=1,\boldsymbol{X}=\boldsymbol{x},R=\textrm{Casecontact tracing},T=1)}{\Pr(I=1|V=0,\boldsymbol{X}=\boldsymbol{x},R=\textrm{Case contact tracing},T=1)}$
$=1-\frac{\Pr(I=1|V=1,\boldsymbol{X}=\boldsymbol{x},R=\textrm{Casecontact tracing})}{\Pr(I=1|V=0,\boldsymbol{X}=\boldsymbol{x},R=\textrm{Case contact tracing})},$ (4)
representing the VE among contacts of known cases.
To connect $VE_{cct}(\boldsymbol{x})$ with the target $VE(\boldsymbol{x})$, we need an additional but reasonable assumption that being a contact of known cases does not modify VE, i.e.,
$\frac{\Pr(I=1|V=1,\boldsymbol{X},R=\textrm{Case contact tracing})}{\Pr(I=1|V=0,\boldsymbol{X},R=\textrm{Case contact tracing})}=\frac{\Pr(I=1|V=1,\boldsymbol{X},R\neq\textrm{Case contact tracing})}{\Pr(I=1|V=0,\boldsymbol{X},R\neq\textrm{Case contact tracing})}.$
This assumption implies $VE_{cct}(\boldsymbol{x})=VE(\boldsymbol{x})$ so that we are able to use the tested case contact subjects to estimate $VE(\boldsymbol{x})$. In practice, close contacts of cases might represent a higher level of exposure intensity to the virus, a potential confounder for VE. Therefore, case contact tracing may cause VE modification if the exposure intensity is not measured, in which cases $VE_{cct}(\boldsymbol{x})\neq VE(\boldsymbol{x})$. On the other hand, whether case contact tracing indicates a higher level of exposure intensity (compared to infection in the general population) needs further justification. This is because some contacts in contact tracing may be casual contacts, and general infections may result from close contacts, perhaps in their household.
Overall, Table 1 summarizes the reasons for testing, their assumptions, and the target VE discussed above. Under minimum assumptions, each category of reasons for testing targets its own VE estimand. In addition, classical TNDs are constrained to healthcare seekers, whereas “disease-unrelated reasons” and “case contact tracing” cover both healthcare-seekers and non-seekers. Categorization of reasons for testing thus demonstrates the bias from an aggregate odds ratio estimator and also provides a stepping stone for strategically combining reasons for testing, which we introduce below.
Table 1: A summary of considered reasons for testing with their assumptions, examples, and target VEs in TNDs. $VE(\boldsymbol{x})$, $VE(\boldsymbol{x},1)$, $VE_{s}(\boldsymbol{x},1)$ and $VE_{cct}(\boldsymbol{x})$ are defined in Equations (1), (2), (3), (4), representing the conditional VE for infection, conditional VE for infection among healthcare-seekers, conditional VE for symptomatic infection (and testing due to symptoms), and conditional VE for infection among contacts of known cases.
Reasons for testing Main assumptions Examples Target VEs
Symptoms (Figure 1) 1. Only healthcare-seekers are tested. 2. Vaccination does not reduce     test-negative symptoms. 3. Given a symptom, the uptake of testing     is the same among cases (test-positives)     and controls (test-negatives). 4. Vaccination has no direct effect     on symptom severity among the infected. Fever, fatigue $VE(\boldsymbol{x},1)$; or $VE_{s}(\boldsymbol{x},1)$ without assumption 4.
Disease- unrelated reasons (Figure 2(a)) 1. Testing is not directly impacted     by infection status or     healthcare-seeking behavior. Surveillance tests, travelling $VE(\boldsymbol{x})$
Case contact tracing (Figure 2(b)) 1. Testing is not directly impacted by     healthcare-seeking behavior. 2. Contact of known cases is not an effect     modifier. Case contact tracing $VE(\boldsymbol{x})$; or $VE_{cct}(\boldsymbol{x})$ without assumption 2.
The above framework also works if the outcome of interest is symptomatic infection, i.e., $I=1$ and $S=1$. With this new outcome definition, testing due to symptoms as in classical TNDs naturally targets $VE_{s}(\boldsymbol{x},1)$ as described in Equation (3). For tests due to disease-unrelated reasons and case contact tracing, the relative risks can be updated by focusing on $I=1,S=1$; this change will require knowing whether tested individuals have symptoms, which may be less realistic in many scenarios. For example, certain prescribed medications may prevent or delay symptom development, testing may block symptoms occurring (particularly if symptoms are severe, e.g., require hospitalization), and symptom evaluation may be unavailable during testing.
Potential bias given various reasons for testing
$\boldsymbol{X}$$H$$V$$I$$S$$R$$T=1$
Figure 3: The DAG for TNDs with multiple testing reasons. Circle nodes stand for random variables, while rectangle nodes represent that the random variables are conditioned on specific values.
When a TND involves multiple reasons for testing (including the above-discussed three categories or other reasons), different causal relationships occur, as demonstrated in Figure 3. For example, when some individuals are tested for onsite work, test results arise from both healthcare-seekers and non-seekers. When testing sometimes occurs due to Covid-specific symptoms (e.g. loss of sense of smell), this may lead to a direct relationship $I\rightarrow T$. Also, in some cases, asymptomatic testing is more likely to occur in healthcare-seekers, leading to a direction relationship $H\rightarrow R$. If such multiple testing reasons all occur in the data, we will obtain a DAG akin to the one shown in Figure 3, where all variables are fully connected. In this case, a naive odds ratio estimator with observed data will be biased for VE, i.e.,
[…]
@@
since the collider bias arising from conditioning on $T=1$ cannot be removed. This bias also exists for estimating $VE(\boldsymbol{x},h)$ for similar reasons and, additionally, due to the unmeasured confounder $H$. As a numerical example, consider a binary confounder $X$ with $\Pr(X=1)=0.5$, $\Pr(V=1|X,H)=0.5$, and $\log\{\Pr(I=1|V,X,H)\}=-0.5-0.5V-1.5X$. We consider testing due to symptoms, screening, or other reasons with
@@
$\Pr(T=1,R=r|I,V,X,H)=\left\{\begin{array}[]{rl}0.2X+0.2IX&\textrm{if}\ r=\textrm{Symptoms},\\ 0.4V&\textrm{if}\ r=\textrm{Screening},\\ 0.1VI&\textrm{if}\ r=\textrm{Other}.\end{array}\right.$
@@
Then, the odds ratio VE estimator based on aggregated results from tested individuals is 0.02, which substantially underestimates the VE among the whole population, which is 0.39. In addition, we have $VE(\boldsymbol{x})=VE(\boldsymbol{x},1)=VE_{s}(\boldsymbol{x},1)=0.39$ in this example, which demonstrates that the aggregated odds ratio estimator may not target any VE of interest.
@@
where $h(V,\boldsymbol{X},r)$ is a user-specified learning model. We can assume $h(V,\boldsymbol{X},r)$ follows a parametric model, such as a log-binomial regression model, or a non-parametric model, using kernel smoothing for non-linear effects of $X$; both approaches are well-developed in the literature with software available for estimation. Compared to parametric models, a non-parametric model requires fewer assumptions and can better approximate the truth with the goal of reducing bias, albeit with the disadvantages of requiring larger sample sizes to avoid overfitting and increased variance associated with a slower convergence rate. We provide mathematical details and model fitting in Web Appendix B. With a user-specified model, we obtain $\widehat{h}(V,\boldsymbol{X},r)$ as the estimate of $h(V,\boldsymbol{X},r)$.
@@
$\widehat{VE}_{s}(\boldsymbol{x},1)$ $=1-\frac{\widehat{h}(V=1,\boldsymbol{X}=\boldsymbol{x},r=\textup{Symptoms})\{1-\widehat{h}(V=0,\boldsymbol{X}=\boldsymbol{x},r=\textup{Symptoms})\}}{\widehat{h}(V=0,\boldsymbol{X}=\boldsymbol{x},r=\textup{Symptoms})\{1-\widehat{h}(V=1,\boldsymbol{X}=\boldsymbol{x},r=\textup{Symptoms})\}},$
$\widehat{VE}(\boldsymbol{x})$ $=1-\frac{\widehat{h}(V=1,\boldsymbol{X}=\boldsymbol{x},r=\textup{Disease-unrelated reasons})}{\widehat{h}(V=0,\boldsymbol{X}=\boldsymbol{x},r=\textup{Disease-unrelated reasons})},$
@@
Under our modeling assumptions and regularity conditions listed in Web Appendix B, the estimators $\widehat{VE}_{s}(\boldsymbol{x},1)$, $\widehat{VE}(\boldsymbol{x})$, and $\widehat{VE}_{cct}(\boldsymbol{x})$ converge in probability to $VE_{s}(\boldsymbol{x},1)$, $VE(\boldsymbol{x})$, and $VE_{cct}(\boldsymbol{x})$ respectively (as the sample size $n$ increases), and their normal-approximated 95% confidence intervals based on estimated SE estimators will have correct coverage in large samples. We also provide variance estimators and technical details in Web Appendix B. If $VE(\boldsymbol{x},1)$ is of interest, an additional assumption is needed to obtain $VE(\boldsymbol{x},1)=VE_{s}(\boldsymbol{x},1)$ (see Table 1), in which case $\widehat{VE}_{s}(\boldsymbol{x},1)$ will also target the estimand $VE(\boldsymbol{x},1)$.
In practice, it is common that the estimands $VE_{s}(\boldsymbol{x},1)$, $VE(\boldsymbol{x},1)$, $VE(\boldsymbol{x})$, and $VE_{cct}(\boldsymbol{x})$ are all distinct. In this case, data collected under disease-unrelated reasons and case contact tracing cannot help with the estimation of $VE_{s}(\boldsymbol{x},1)$ or $VE(\boldsymbol{x},1)$. However, the additional data are not useless: they can be used to estimate $VE(\boldsymbol{x})$, and $VE_{cct}(\boldsymbol{x})$; both estimands can be of interest depending on the nature of vaccination (intervention) and the scientific question.
@@
In special cases where several VE estimands coincide, we are able to combine the corresponding estimators to yield a more precise estimator. If there is evidence that case contact tracing does not modify VE, we have $VE_{cct}(\boldsymbol{x})=VE(\boldsymbol{x})$ and propose a combined estimator for $VE(\boldsymbol{x})$, defined as
$\widehat{VE}_{comb}(\boldsymbol{x})$ $=w_{1}\widehat{VE}(\boldsymbol{x})+w_{2}\widehat{VE}_{cct}(\boldsymbol{x}),$
@@
If we assume healthcare-seeking is not an effect modifier, vaccination has no direct effect on symptom alleviation, and case tracing does not modify VE, then $VE_{s}(\boldsymbol{x},1)=VE(\boldsymbol{x})=VE_{cct}(\boldsymbol{x})$, and we can construct an estimator combining all three categories of reasons:
$\widehat{VE}^{*}_{comb}(\boldsymbol{x})$ $=w_{1}^{*}\widehat{VE}_{s}(\boldsymbol{x},1)+w_{2}^{*}\widehat{VE}(\boldsymbol{x})+w_{3}^{*}\widehat{VE}_{cct}(\boldsymbol{x})$
where $(w_{1}^{*},w_{2}^{*},w_{3}^{*})$ are arbitrary positive weights with $w_{1}^{*}+w_{2}^{*}+w_{3}^{*}=1$, and these weights can be chosen similarly to $(w_{1},w_{2})$. When their assumptions hold, both combined estimators target $VE(\boldsymbol{x})$, and $\widehat{VE}^{*}_{comb}(\boldsymbol{x})$ improves precision over $\widehat{VE}_{s}(\boldsymbol{x},1)$ by leveraging the additional data. The amount of precision improvement will be greater if we have more data on asymptomatic tests. In the Covid-19 context, asymptomatic tests can often account for a large proportion of data, thereby having the potential to improve precision substantially. This result extends [13], which proposed combining stratified estimators for symptoms and case contact tracing (without specifying how to choose the weights) but did not consider disease-unrelated reasons or explain the underlying assumptions for estimation validity.
In fact, whether the above VE estimands are equal is testable from data. To evaluate whether case contact tracing modifies the VE, we can test the hypothesis $H_{0}:VE_{cct}(\boldsymbol{x})=VE(\boldsymbol{x})$ for any $\boldsymbol{x}$ such that $\widehat{VE}_{cct}(\boldsymbol{x})$ and $\widehat{VE}(\boldsymbol{x})$ can be constructed. Since both $\widehat{VE}_{cct}(\boldsymbol{x})$ and $\widehat{VE}(\boldsymbol{x})$ are normally distributed (on the root-$n$ scale) and independent in large samples, $\widehat{VE}_{cct}(\boldsymbol{x})-\widehat{VE}(\boldsymbol{x})$ also follows a normal distribution in large samples, from which $H_{0}$ can be rejected if $|\widehat{VE}_{cct}(\boldsymbol{x})-\widehat{VE}(\boldsymbol{x})|$ exceeds the critical value for a given type I error. Following a similar logic, we can also test $H_{0}:VE_{s}(\boldsymbol{x},1)=VE(\boldsymbol{x})$ using $\widehat{VE}_{s}(\boldsymbol{x},1)$ and $\widehat{VE}(\boldsymbol{x})$. This test has greater power if the candidate VEs have substantial differences, the sample sizes are large, and the VE estimates have small variances.
@@
In Web Appendix C, we provide simulation studies to demonstrate our results numerically. In summary, the simulation studies show that (1) the traditional aggregated odds ratio estimator can be biased for TNDs with multiple reasons for testing, (2) the proposed estimators are unbiased for the target estimands (with the necessary assumptions), (3) leveraging data for asymptomatic reasons for testing can improve precision. Furthermore, supplementary simulations to explore the effect of misclassification of the infection status and missing values in the reasons for testing are also provided in Web Appendix C.

### v3

@@ Introduction
since the collider bias arising from conditioning on $T=1$ cannot be removed. This bias also exists for estimating $VE_{symp}(\boldsymbol{x})$ or $VE_{symp}(\boldsymbol{x},1)$ for similar reasons and, additionally, due to the unmeasured confounder $H$. As a numerical example, consider $\Pr(H=1)=0.5$, $\Pr(V=1|H)=0.5$, $\log\{\Pr(I=1|V,H)\}=-0.5-0.5V-1.5H$, and $\Pr(S=1|V,H,I)=0.5$. We consider testing due to symptoms, mandatory screening, or other reasons with
@@
$\Pr(R=r,T=1|\mathcal{Z})=\left\{\begin{array}[]{rl}0.2HS&\textrm{if}\ r=\textrm{Symptoms},\\ 0.2V+0.1&\textrm{if}\ r=\textrm{Mandatory screening},\\ 0.3VI&\textrm{if}\ r=\textrm{Others}.\end{array}\right.$
@@
Then, the odds ratio estimator based on aggregated results from tested individuals is $-0.02$, which substantially underestimates the VE among the whole population, which is 0.39. In addition, we have $VE(\boldsymbol{x})=VE_{symp}(\boldsymbol{x})=VE_{symp}(\boldsymbol{x},1)=0.39$ in this example, which demonstrates that the aggregated odds ratio estimator may not target any VE of interest.
Estimating the VE parameters
@@
where $h(V,\boldsymbol{X},r)$ is a user-specified learning model. We can assume $h(V,\boldsymbol{X},r)$ follows a parametric model, such as a log-binomial regression model, or a non-parametric model, using kernel smoothing for non-linear effects of $X$; both approaches are well-developed in the literature with software available for estimation. Compared to parametric models, a non-parametric model requires fewer assumptions and can better approximate the truth with the goal of reducing bias, albeit with the disadvantages of requiring larger sample sizes to avoid overfitting and increased variance associated with a slower convergence rate. We provide mathematical details and model fitting in Web Appendix C. With a user-specified model, we obtain $\widehat{h}(V,\boldsymbol{X},r)$ as the estimate of $h(V,\boldsymbol{X},r)$.
@@
$\widehat{VE}_{symp}(\boldsymbol{x},1)$ $=1-\frac{\widehat{h}(V=1,\boldsymbol{X}=\boldsymbol{x},r=\textup{Symptoms})\{1-\widehat{h}(V=0,\boldsymbol{X}=\boldsymbol{x},r=\textup{Symptoms})\}}{\widehat{h}(V=0,\boldsymbol{X}=\boldsymbol{x},r=\textup{Symptoms})\{1-\widehat{h}(V=1,\boldsymbol{X}=\boldsymbol{x},r=\textup{Symptoms})\}},$
$\widehat{VE}(\boldsymbol{x})$ $=1-\frac{\widehat{h}(V=1,\boldsymbol{X}=\boldsymbol{x},r=\textup{Mandatory screening})}{\widehat{h}(V=0,\boldsymbol{X}=\boldsymbol{x},r=\textup{Mandatory screening})},$
@@
The estimators $\widehat{VE}_{symp}(\boldsymbol{x})$ and $\widehat{VE}_{symp,cct}(\boldsymbol{x})$ can be constructed similarly by replacing $I=1$ by $I=1,S=1$ in the definition of $h$, using the data from $r=\textup{Mandatory screening}$ and $r=\textup{Case contact tracing}$, respectively.
Under our modeling assumptions and regularity conditions listed in Web Appendix C, each estimator converges in probability to their targe VE parameter as the sample size $n$ increases, and their normal-approximated 95% confidence intervals based on estimated standard error estimators will have correct coverage in large samples. We also provide variance estimators and technical details in Web Appendix C.
Connecting VE parameters
In practice, it is common that the above-discussed VE parameters are all distinct. However, with additional assumptions, we are able to connect different VE parameters. For testing due to symptoms, if VE does not vary by the healthcare-seeking behavior [1] and $V$ is independent of $H$ given $\boldsymbol{X}$, then $VE_{symp}(\boldsymbol{x},1)$ will be equal to $VE_{symp}(\boldsymbol{x})$, which motivates us to combine the VE estimation from different testing reasons. For case contact tracing, we will have $VE_{cct}(\boldsymbol{x})=VE(\boldsymbol{x})$ and $VE_{symp,cct}(\boldsymbol{x})=VE_{symp}(\boldsymbol{x})$, if being a contact of known cases does not modify VE and $V$ is independent of $C$ given $\boldsymbol{X}$. Nevertheless, this assumption needs further justification since close contact of cases might represent a higher level of exposure intensity to the virus, a potential confounder for VE.
For the VEs we have discussed, it is important to highlight that the VE against symptomatic infection and VE against general infections are different. Since vaccination may further alleviate symptoms beyond guarding against infection, $VE_{symp}(\boldsymbol{x})$ is generally larger than $VE(\boldsymbol{x})$. Therefore, while data from different sources can be combined for a joint analysis, $VE(\boldsymbol{x})$ and $VE_{symp}(\boldsymbol{x})$ are estimated separately below.
In fact, whether the above VE parameters are equal is testable from the data. To evaluate whether case contact tracing modifies the VE, we can test the hypothesis $H_{0}:VE_{cct}(\boldsymbol{x})=VE(\boldsymbol{x})$ for any $\boldsymbol{x}$ such that $\widehat{VE}_{cct}(\boldsymbol{x})$ and $\widehat{VE}(\boldsymbol{x})$ can be constructed. Since both $\widehat{VE}_{cct}(\boldsymbol{x})$ and $\widehat{VE}(\boldsymbol{x})$ are normally distributed (on the root-$n$ scale) and independent in large samples, $\widehat{VE}_{cct}(\boldsymbol{x})-\widehat{VE}(\boldsymbol{x})$ also follows a normal distribution in large samples, from which $H_{0}$ can be rejected if $|\widehat{VE}_{cct}(\boldsymbol{x})-\widehat{VE}(\boldsymbol{x})|$ exceeds the critical value for a given type I error. Following a similar logic, we can also test $H_{0}:VE_{symp}(\boldsymbol{x},1)=VE_{symp}(\boldsymbol{x})$ using $\widehat{VE}_{symp}(\boldsymbol{x},1)$ and $\widehat{VE}_{symp}(\boldsymbol{x})$. This test has greater power if the candidate VEs have substantial differences, the sample sizes are large, and the VE estimates have small variances.
@@
In special cases where several VE estimands coincide, we are able to combine the corresponding estimators to yield a more precise estimator. If $VE_{cct}(\boldsymbol{x})=VE(\boldsymbol{x})$, we propose a combined estimator for $VE(\boldsymbol{x})$, defined as
$\widehat{VE}^{*}(\boldsymbol{x})$ $=w_{1}\widehat{VE}(\boldsymbol{x})+w_{2}\widehat{VE}_{cct}(\boldsymbol{x}),$
@@
If $VE_{symp}(\boldsymbol{x},1)=VE_{symp}(\boldsymbol{x})=VE_{symp,cct}(\boldsymbol{x})$, we can construct an estimator combining all three categories of reasons:
$\widehat{VE}^{*}_{symp}(\boldsymbol{x})$ $=w_{1}^{*}\widehat{VE}_{symp}(\boldsymbol{x},1)+w_{2}^{*}\widehat{VE}_{symp}(\boldsymbol{x})+w_{3}^{*}\widehat{VE}_{symp,cct}(\boldsymbol{x})$
where $(w_{1}^{*},w_{2}^{*},w_{3}^{*})$ are arbitrary positive weights with $w_{1}^{*}+w_{2}^{*}+w_{3}^{*}=1$, and these weights can be chosen similarly to $(w_{1},w_{2})$. When their assumptions hold, combined estimators improve precision by leveraging the additional data. The amount of precision improvement will be greater if we have more data on asymptomatic tests. In the Covid-19 context, asymptomatic tests can often account for a large proportion of data, thereby having the potential to improve precision substantially. This result extends [13], which proposed combining stratified estimators for symptoms and case contact tracing (without specifying how to choose the weights) but did not consider mandatory screening or explain the underlying assumptions for estimation validity.
Numerical illustration
@@
In Web Appendix D, we provide simulation studies to demonstrate our results numerically. In summary, the simulation studies show that (1) the traditional aggregated odds ratio estimator can be biased for TNDs with multiple reasons for testing, (2) the proposed estimators are unbiased for the target estimands (with the necessary assumptions), (3) leveraging data for asymptomatic reasons for testing can improve precision. Furthermore, supplementary simulations to explore the effect of misclassification of the infection status and missing values in the reasons for testing are also provided in Web Appendix D.
Data application
We use the electronic health records (EHR) within the University of Michigan Health System for a realistic demonstration of our results. Focusing on the period between July 1 and December 31, 2021 (approximately the Delta outbreak), we extract the vaccination history, laboratory results of SARS-CoV-2 tests (positive/negative), Covid-like symptom evaluation (yes/no), and covariates including gender, body mass index, and age. We let $V=1$ if a subject received any vaccine dose. Since reasons for testing are not explicitly recorded in the EHR, we define testing due to symptoms as outpatient testing on symptomatic subjects, and case contact tracing as those with ICD-10 code Z20.822, meaning “Contact with and (suspected) exposure to Covid-19”. For mandatory screening, while the ICD-10 code Z11.52 matches “screening for Covid-19”, it rarely appeared in the EHR; we therefore exclude this reason for testing from our analysis. It is important to highlight that our definitions of reasons for testing are only for demonstration, which may deviate from the truth and fail in other datasets.
A parametric generalized linear model is used to estimate $VE_{symp}(\boldsymbol{x},1)$ based on testing due to symptoms and $VE_{symp,cct}(\boldsymbol{x})$ based on case contact tracing. First, we implement the regression model without vaccination-covariate interactions, which results in constant VE estimates across strata. Table 2 summarizes the analysis results and shows a significant difference in VEs between the two categories (p-value $<0.001$), which may reflect the effect modification from healthcare-seeking behavior and contact with cases. As a result, we do not perform the combination of the stratified estimators. However, we implemented the odds ratio estimator aggregating both reasons and obtained $40\%$ with standard error $0.04$. This estimator shows bias to either VE, which further demonstrates the issue of an unstratified analysis. Furthermore, Figure 5 shows how the VE changes with age through a vaccination-age interaction term in the regression, and the results indicate better VE for older individuals in both categories.
Table 2: Summary results of data application.
Reason Sample size VE Standard error
Symptoms 1087 0.77 0.05
Case contact tracing 22106 0.33 0.05
Figure 5: VE by age in the data application. The red and blue colors represent tests due to symptoms and case contact racing, respectively.